OpenStudy (anonymous):

Hey, can you guys help me with studying? I'm studying for a test tonight. Yes, because I study online and I have to take a geometry test, so I want to be prepared. I am having problems with this unit but I have problems with math everyday lol so the only way to get decent grades is for me to try to understand .so please help?

OpenStudy (pphalke):

i might be able to

OpenStudy (anonymous):

It is geometry, is that alright with you?

OpenStudy (fortytherapper):

Let's see it

OpenStudy (anonymous):

Okay! Well here is one, let me just attach it in a minute

OpenStudy (anonymous):

It confuses me, aren't they similar triangles?

OpenStudy (anonymous):

I set up a proportion though, I think it might be right. 6.3/11.2 = 5.4/x

563blackghost (563blackghost):

That would be correct ^^ just cross multiply to find x

OpenStudy (anonymous):

Hmm, okay so I got 9.6? Would that be right?

563blackghost (563blackghost):

That would be correct ^^

OpenStudy (anonymous):

This is a bit harder, because I don't exactly know how to find values like this, and it seems to me I cannot use a proportion in this case :/

563blackghost (563blackghost):

one sec...

OpenStudy (anonymous):

Okay! :)

OpenStudy (fortytherapper):

Looks like an Isosceles Triangle (The small triangle) because it looks to be 45-45-90

OpenStudy (anonymous):

Yeah, but what can I do with that? Does that allow me to set another proportion? Lol I'll solve anything with a proportion if I'm allowed.

OpenStudy (fortytherapper):

Since it's both Isosceles and a right triangle (It doesn't say it, but I assume) that means the two sides that aren't the hypotenuse are equal

OpenStudy (anonymous):

Hmm. the thing is, usually the hypotenuse is the longest side, so would it say, that it would be 29/21?

OpenStudy (mathmale):

To avoid confusion, please post just one problem at a time, not several. And in the first problem, no, the two triangles are not similar. Please share what you have done towards solving each problem before asking for help. Thx.

rishavraj (rishavraj):

dont think because of the look... :P see there are two riight angled triangles ...apply pythogoreas theorem

OpenStudy (anonymous):

The answer choices ,which I do have a few, they don't make sense to me. Here is what they say: a) a = 400/21 , b =580 / 21 (where are they getting these huge numbers?) b) a = 400/21, b = 20/21 c) a = 580 / 21, b = 29/21 and d) 20/21, b = 580 / 21

rishavraj (rishavraj):

@563blackghost u got the solution ....coz I am not gonna solve it rn -_-

OpenStudy (anonymous):

Oh no, you guys, I don't which for you guys to solve it, I just need help doing it myself hahaha

563blackghost (563blackghost):

:P im trying to think of what collary would be used to solve this @rishavraj

563blackghost (563blackghost):

Look at this....

OpenStudy (anonymous):

I have this on my notes: hypotenuse / leg = leg/ adjacent segment to the leg

563blackghost (563blackghost):

From the smaller inner triangle we see that 202+a2=b2 This is by the Pythagorean Theorem. We also see form the larger outer triangle 292+b2=(21+a)2 Also by the Pythagorean Theorem. We now have a system of equations to work with. Substituting 202+a2 in for b2 in the second equation we have: 292+(202+a2)=(21+a)2 Simplifying the left and right hand sides, we have: 841+400+a2=441+a2+42a 1241+a2=441+a2+42a now we cancel the (a^2)s and solve for a 1241+a2=441+a2+42a 800=42a ~ @Ephilo ~ http://openstudy.com/updates/51239016e4b02ca4087abaa4

563blackghost (563blackghost):

so @rishavraj was correct

OpenStudy (anonymous):

Hmm, I'm not very familiar with that Theorem. We've only gone over it super briefly.

OpenStudy (anonymous):

Also, I apologize for posting more than question. I hadn't seen that message (things don't load fast on my computer)

OpenStudy (nikato):

actually, its another proportion problem. just needs a little more set up before beginning to solve

OpenStudy (anonymous):

I think so. I feel like I need to use one of the corollaries that @563blackghost gave me. I just don't know where to place the values, and it is on my notes and I'm looking at it, and it doesn't make sense to me, but I know I have to use one of the two.

OpenStudy (nikato):

this is how i would do it.  so we are given this. and you should know since its a straight line, the other angle should also be a right angle. and you're given two angles are similar, so you can use AA to prove the 2 triangles are similar

OpenStudy (nikato):

now draw them side by side  now you can set up proportions to solve for a and b

OpenStudy (anonymous):

Yes I can! Thank you hahaha sorry that it took me long, internet in my hosue is not as reliable

OpenStudy (nikato):

haha. no problem. hope it helps you on your test

OpenStudy (anonymous):

Well, I'm almost done with my study guide. My main problem is figuring out things like this one. I have like four of them left, and all of them are somewhat like this but more comp[plicated so I really appreciate it

OpenStudy (nikato):

i see. well if you need more help, you should probably close this one and open annother one